OpenStudy (luigi0210):

I need a bit of help, Exact Equations:

OpenStudy (luigi0210):

\(\LARGE (tanx-sinx ~siny)dx+cosx~cosy~dy=0 \) I don't understand how to find the partials

OpenStudy (anonymous):

intrtesting

OpenStudy (alekos):

You trying to solve for y?

OpenStudy (valpey):

\[\frac{dy}{dx} = \frac{-\tan x+\sin x \sin y}{\cos x \cos y}\] then integrate while treating sin y and cos y as constants, then repeat for dx/dy (I think)

OpenStudy (alekos):

Integrate the entire expression 1st part wr to x and 2nd part wr to y

OpenStudy (freckles):

\[M=\tan(x)-\sin(x)\sin(y) \\ N=\cos(x)\cos(y) \\ M_y=(\tan(x)-\sin(x)\sin(y))_y=(\tan(x))_y-(\sin(x)\sin(y))_y \\ M_y=0-\sin(x)(\sin(y))_y =-\sin(x)(\cos(y))\] \[\text{ Now you need to find } N_x \]